OpenStudy (anonymous):

What is the inverse of the function below? f(x) = 11x + 9

OpenStudy (anonymous):

If f(x) sends x to y, i.e.\[y=11x+9\] then the inverse function sends y to x, i.e.\[x=11y+9\] so if you solve for y in the second equation, replace f(x) with f^-1(x) for the inverse function.

OpenStudy (anonymous):

A. f -1(x) = -11x - 9 B. f -1(x) = x - 9 11 C. f -1(x) = x - 11 9 D. f -1(x) = 9 + 11x @xeph

OpenStudy (anonymous):

nvm